OpenStudy (anonymous):

For how many positive integers \(n\)\[\left \lfloor \frac{n^2}{3} \right \rfloor\]is a prime number? \(\left \lfloor \right \rfloor\) denotes the floor function.

ganeshie8 (ganeshie8):

my initial guess is to use little fermat for \((n,3)=1\) we have \[n^2\equiv 1\pmod{3}\]

OpenStudy (anonymous):

what comes after that?

ganeshie8 (ganeshie8):

\[\left \lfloor \frac{n^2}{3} \right \rfloor = \left \lfloor \frac{3k+1}{3} \right \rfloor = k\] ?

OpenStudy (anonymous):

well, that's right, but how do we know that \(k\) is a prime or not?

OpenStudy (anonymous):

Use elementary methods. Just simplify the expression.

ganeshie8 (ganeshie8):

\[n^2 = 3k+1 \implies 3k = n^2-1=(n-1)(n+1)\]

OpenStudy (p0sitr0n):

For a general field \[\mathbb{F}=\mathbb{R}\] this simplification won't work tho , since the floor function has an input interval that equals all the same value, i.e. \[\lfloor5.4\rfloor=\lfloor5.2\rfloor \nrightarrow 5.4 = 5.2\], so careful

ganeshie8 (ganeshie8):

that yields \(\large k\in \{3,5\}\) consequently \(n\in \{4\}\)

ganeshie8 (ganeshie8):

for \((n,3)\ne 1\), \(\left \lfloor \frac{n^2}{3} \right \rfloor =\left \lfloor \frac{(3k)^2}{3} \right \rfloor = 3k^2\) is always composite except for \(k=1\)

ganeshie8 (ganeshie8):

Overall \(n\in\{3,4\}\)

OpenStudy (anonymous):

how about \(n=3k+2\)

OpenStudy (anonymous):

and \(n=3k+1\)?

OpenStudy (anonymous):

sry for late response, I was out

ganeshie8 (ganeshie8):

little fermat covers both n=3k+1, 3k+2 right

OpenStudy (anonymous):

right :)

ganeshie8 (ganeshie8):

would love to see an alternate method as my method above is pretty hacky

ganeshie8 (ganeshie8):

another way to look at it is : \(n^2\) can never be \(3k+2 \) because \((3k\pm 1)^2 = 3M+1\). so \(n^2\equiv 0,1\pmod{3}\)